OpenStudy (anonymous):

Graph the two equations to solve the system. What is the solution? 3x - 2y = 6 and 5x + 3y = -9 1) (0,0) 2) (0,-3) 3) No Solution 4) Infinity solutions.

OpenStudy (candyme):

To graph these, first get the two equations into y=mx+b form.

OpenStudy (candyme):

3x-2y=6 simplifies to y=3/2x-3 5x+3y=-9 simplifies to y=-5/3x-3